OpenStudy (anonymous):

A mother wants to invest $11,000 on her childs education, she invests a portion of the money in a bank certificate of deposit (CD account) which earns 4% and the rest in a savings bond that earns 7% If the total interest earned is $660, how much was invested at each rate?

OpenStudy (anonymous):

660=.04a+(11000-a)*.07

OpenStudy (anonymous):

a+b=11,000..a and b are the two parts of the 11000 a=11000-b...solve for a .04a+.07b=660....formula for the interest total .04(11000-b)+.07b=660...plug in a and solve for b, then refer back to first equation to get a

OpenStudy (anonymous):

660=-.03a+770 110=.03a

OpenStudy (anonymous):

What is the Amount for the CD account, the 110=3666.68 right?